OpenStudy (anonymous):

HELP PLEASE! How do you graph (y-2)^2/4 - (x+3)^2/9 =1

OpenStudy (ajanijones):

1 sec let me do the work.

OpenStudy (anonymous):

\[\frac{ (y-2)^{2} }{ 4} - \frac{ (x+3) ^{2}}{ 9 } =1\]

OpenStudy (anonymous):

Okay :)

OpenStudy (ajanijones):

sorry i have to go sorry

OpenStudy (anonymous):

;/